15. A stock has a market price of $43.45 and pays a $1.75 dividend. What is the dividend yield?
a. 3.12 percent
b. 4.03 percent
c. 2.38 percent
d. 1.94 percent

16. Willow Co. pays an annual dividend of $1.45 per share and sells for $25.50 a share based on a market rate of return of 12.5 percent. What is the capital gain yield?
a. 4.41 percent
b. 7.17 percent
c. 5.12 percent
d. 6.81 percent
Answer #1

a.dividend yield=Annual dividend/Current value

=1.75/43.45

= 4.03 percent%(Approx).

b.dividend yield=Annual dividend/Current value

=(1.45/25.5)=5.69%(Approx)

capital gain yield=Market rate of return-Dividend yield

=12.5-5.69

=6.81 percent(Approx).